The energy of electron in first energy level is -21.79*10-12erg per atom. What would be the energy of electron in second energy level?
1
Expert's answer
2018-07-23T06:43:22-0400
The total energy, E of the electron is the sum of kinetic energy and potential energy.

Kinetic energy of the electron = ½ mv2

Potential energy = -KZe2 /r

Total energy = 1/2 mv2 – KZe2 /r

Total energy (E) = KZe2 /2r – KZe2 /r = -KZe2 /2r

Substituting for r, gives us

E = 2π2mZ2e'K2/n2h2 where n = 1, 2, 3………

Only certain energies are allowed to the electron. This energy expression consist of many fundamental constant.

The simplified expression:

E = –21.79x10–12 x z2n2 erg per atom

The energy of electron in second energy level would be:

E = –87.16x10–12 erg per atom.